### v1

# Symmetry Energy Expansion with Strange Dense Matter

Abstract
The Quantum Chromodynamics phase diagram at large densities and low temperature can be probed both by neutron stars and low-energy heavy-ion collisions. Heavy-ion collisions are nearly isospin symmetric systems whereas neutron stars are highly isospin asymmetric since they are neutron-rich. The symmetry energy expansion is used to connect these regimes across isospin asymmetry. However, the current symmetry energy expansion does not account for strange particles. Here we include a finite strangeness to calculate isospin asymmetry and find that the symmetry energy expansion has a skewness term in the presence of strangeness for the case of weak equilibrium.
Introduction Depending on the temperatures $T$, baryon densities $n_{B}$, electric charge densities $n_{Q}$, and strangeness densities $n_{S}$ reached, nuclear matter may have drastically different features Collins and Perry (1975); Cabibbo and Parisi (1975); Bzdak et al. (2020); An et al. (2022); Sorensen et al. (2024); Lovato et al. (2022). In the lab, one generally discusses the charge fraction $Y_{Q}=Z/A=n_{Q}/n_{B}$ because it is dictated by the choice of nuclei with proton number $Z$ and total number of nucleons $A$. Similarly, one can define a strangeness fraction of $Y_{S}=n_{S}/n_{B}$. The Quantum Chromodynamics (QCD) phase diagram at finite density is a 4-dimension $\left\{T,n_{B},Y_{Q},Y_{S}\right\}$ space that can be probed by lab experiments (heavy-ion collisions Danielewicz et al. (2002); Pratt et al. (2015); Ablyazimov et al. (2017); Alba et al. (2018); Spieles and Bleicher (2020); Monnai et al. (2021); Oliinychenko et al. (2023); Yao et al. (2024); Plumberg et al. (2024)), astrophysical objects (neutron stars Abbott et al. (2017); Miller et al. (2019, 2021); Riley et al. (2019, 2021)), or astrophysical events (binary neutron star mergers Most et al. (2019), supernovae Zha et al. (2020), or even the Big Bang Borsanyi et al. (2016)).
In the low $T$ and large $n_{B}$ limit, we have three different systems to probe the phase diagram: i) neutron stars at low $T\sim 0$ with small $Y_{Q}\lesssim 0.1$; ii) neutron star mergers at $T\sim[0,100]\,\rm{MeV}$ with $Y_{Q}\sim[0.01,0.2]$; iii) low-energy heavy-ion collisions at $T\sim[50,150]\,\rm{MeV}$ with $Y_{Q}\sim[0.38,0.5]$. Values of $Y_{S}$ are often excluded from discussions of the phase diagram because heavy-ion collisions probe global $Y_{S}=0$ (due to strangeness conservation in QCD) Monnai et al. (2019); Noronha-Hostler et al. (2019); Karthein et al. (2021), [^1] This does not exclude local fluctuations in $Y_{S}$ Xu (2017); Noronha-Hostler et al. (2016); Bluhm and Nahrgang (2019); Alba et al. (2017); Bellwied et al. (2019, 2020); Alba et al. (2020); San Martin et al. (2023). and because, even though an average $Y_{S}<0$ is allowed by weak interactions in neutron stars, the presence of strangeness in neutron star matter is not settled. [^2] If strangeness is present, then $Y_{S}<0$ is expected, since strange quarks carry negative strangeness and the strange antiquark is suppressed due to its negative baryon number. Confirming the presence of strange particles within the core of neutron stars is of great importance to the nuclear physics community Kaplan and Nelson (1986); Glendenning (1992); Schaffner et al. (1994, 1993); Hatsuda and Kunihiro (1994); Pisarski (2000); Buballa (2005); Weber (2005); Weissenborn et al. (2012); Dexheimer et al. (2015); Gomes et al. (2015); Tolos and Fabbietti (2020), as it would establish the first known regime with a stable $Y_{S}\neq 0$ in the universe. In fact, phase transitions can become significantly more complicated when multiple conserved charges are present Glendenning (1992). However, determining the properties at the core of a neutron star is not yet possible with current astrophysical data Mroczek et al. (2024).
Thus, forging phenomenological tools to connect neutron stars to heavy-ion collisions is necessary to provide further constraints to the neutron star equation of state (EOS) to determine its interior structure (see e.g. Yao et al. (2024)). Already, the symmetry energy expansion Li et al. (1998); Chen et al. (2005); Li et al. (2019); Xu et al. (2010); Li et al. (2021) can connect the limits of isospin symmetric nuclear matter (SNM) $Y_{Q}=0.5$ and pure neutron matter (PNM) $Y_{Q}=0$ to isospin asymmetric nuclear matter (ANM) found in neutron stars, where $Y_{Q}(n_{B})$ has a functional form that depends on $n_{B}$, determined by $\beta$-equilibrium. The symmetry energy expansion has been used both to connect estimates for the EOS in heavy-ion collisions and in neutron stars and to extend theoretical results (such as the EOS from chiral effective field theory Wellenhofer et al. (2015, 2016); Drischler et al. (2020, 2021)) from the SNM/PNM limits to the case of neutron star matter at $\beta$-equilibrium. Only recently Yao et al. (2024), this expansion has been used to convert a family of neutron star EOS to the SNM regime and to then test them against heavy-ion flow data (while eliminating many EOS through causality, stability, and saturation property constraints), opening up a new path to further constrain the neutron star EOS.
However, the usual symmetry energy expansion cannot handle $Y_{S}\neq 0$, such that it cannot be used with current constraints from heavy ions or neutron stars to test the possibility of strange particles within neutron stars. Some groups have attempted to add strangeness considerations to the expansion Bedaque and Steiner (2015) or to calculate the influence of strangeness on the symmetry energy coefficients Providencia and Rabhi (2013); Providência et al. (2014); Bednarek et al. (2019) but did not consider the influence of strangeness on isospin asymmetry. That being said, there is a straightforward method to incorporate $Y_{S}\neq 0$ into the symmetry energy expansion based on its contribution to isospin, letting us take advantage of the approximate isospin symmetry, which we do for the first time in the work [^3] Previous work did study this within a hadron resonance gas Mekjian (2007a, b) but not on the symmetry energy expansion. We consider two cases: an isospin symmetric basis vs the weak equilibrium case where the strangeness chemical potential vanishes $\mu_{S}=0$. We then show that our method can reproduce a chiral mean field (CMF) model Cruz-Camacho et al. (2024) at $T=0$ as one changes the $Y_{Q}$. We discuss new theoretical and experimental methods to put constraints on the expansion coefficients in the strange symmetry energy expansion.
Notation: We define chemical potentials in the basis of baryon number $B$, charge $Q$, and strangeness $\mathcal{S}$ as [^4] Note that $\mu_{B}$ and $\mu_{S}$ become different if isospin is fixed instead of charge. See Ref. Aryal et al. (2020) and the Supplemental Material for details.
$\mu_{B}\equiv\left(\frac{\partial E}{\partial B}\right)_{Q,\mathcal{S}},$ $\mu_{Q}\equiv\left(\frac{\partial E}{\partial Q}\right)_{B,\mathcal{S}},$ $\mu_{S}\equiv\left(\frac{\partial E}{\partial\mathcal{S}}\right)_{B,Q}$ (1)
where $E$ is the internal energy of nuclear matter. For a given particle $i$ an effective chemical potential can be defined such that
$\tilde{\mu}_{i}=B_{i}\mu_{B}+S_{i}\mu_{S}+Q_{i}\mu_{Q}.$ (2)
The isospin along the neutron-proton direction can be related to the conventional $(B,Q,\mathcal{S})$ basis through the Gell-Mann-Nishijima formula Nakano and Nishijima (1953); Gell-Mann (1956) [^5] In principle, other flavors could be added but we do not expect charm or any heavy quarks to appear within neutron stars.,
$Q=I_{z}+\frac{1}{2}\left(B+\mathcal{S}\right),$ (3)
which can be used to rewrite Eq. (2) in the $(B,I_{z},\mathcal{S})$ basis,
$\tilde{\mu}_{i}$ $=B_{i}\left(\mu_{B}+\frac{1}{2}\mu_{Q}\right)+S_{i}\left(\mu_{S}+\frac{1}{2}\mu_{Q}\right)+I_{z,i}\mu_{Q}$ (4)
$=B_{i}\mu_{B}^{(S,I_{z})}+S_{i}\mu_{S}^{(B,I_{z})}+I_{z,i}\mu_{I},$
where
$\mu_{B}^{(S,I_{z})}$ $\equiv\left(\frac{\partial E}{\partial\mathcal{B}}\right)_{S,I_{z}}=\mu_{B}+\frac{1}{2}\mu_{Q},$ (5a)
$\mu_{S}^{(B,I_{z})}$ $\equiv\left(\frac{\partial E}{\partial\mathcal{S}}\right)_{B,I_{z}}=\mu_{S}+\frac{1}{2}\mu_{Q}$ (5b)
$\mu_{I}$ $\equiv\left(\frac{\partial E}{\partial I_{z}}\right)_{B,\mathcal{S}}=\mu_{Q},$ (5c)
such that one can choose to write the chemical potentials either in the $BSQ$ basis of $\left\{\mu_{B},\mu_{S},\mu_{Q}\right\}$ or the $BSI_{z}$ basis of $\left\{\mu_{B}^{(S,I_{z})},\mu_{S}^{(B,I_{z})},\mu_{I}\right\}$.
Isospin Asymmetry with $\bm{Y_{S}\neq 0}$ In a system with more than one conserved charge, charge conjugation symmetry requires that all charges be reflected. In the context of QCD, this yields $E(B,Q,\mathcal{S})=E(-B,-Q,-\mathcal{S})$, which is not very helpful when we are interested in large positive values of the baryon number only. [^6] This symmetry becomes relevant for lower values of the baryon density, such as can be found in heavy-ion collisions. However, thanks to the approximate isospin symmetry of QCD, another relationship can be found: $E(B,I_{z},\mathcal{S})\approx E(B,-I_{z},\mathcal{S})$, which holds to very good approximation. This approximate isospin flipping $\mathbb{Z}_{2}$ symmetry of $E$ is more relevant for large values of $B$ and simplifies the symmetry energy expansion.
Similar to how Eq. (3) relates the charges, one can divide Eq. (3) by $B$ to obtain a relation between charge fractions Aryal et al. (2021)
$Y_{I}=Y_{Q}-\frac{1}{2}-\frac{1}{2}Y_{\mathcal{S}}.$ (6)
In the absence of strangeness, $Y_{S}=0$, isospin reversal symmetry $I_{z}\to-I_{z}$ is equivalent to
$\delta_{Q}\to-\delta_{Q},$ $\delta_{Q}\equiv 1-2\,Y_{Q},$ $(Y_{S}=0)$ (7)
which is the traditional isospin asymmetry parameter. This reflection symmetry is related to the absence of odd terms in $\delta_{Q}$ in the usual symmetry energy expansion, but fails for $Y_{S}\neq 0$. Thus, in the presence of strangeness, we have instead the reflection symmetry
$\delta_{I}\to-\delta_{I},$ $\delta_{I}\equiv 1+Y_{S}-2\,Y_{Q}=-2\,Y_{I},$ $(Y_{S}\neq 0)$ (8)
which generalizes Eq. (7).
The inclusion of $Y_{S}$ in Eq. (8) is necessary for two reasons: i) the isospin reversal $I_{z}\to-I_{z}$ operation does not commute with the change of basis $(B,I_{z},\mathcal{S})\to(B,Q,\mathcal{S})$ (see the Supplemental Material), and ii) unlike isospin, there is no (approximate) reversal symmetry $\mathcal{S}\to-\mathcal{S}$, unless one also takes $B\to-B$, such that the reflection symmetry under $\delta\to-\delta$ is lost if $Y_{S}$ is not included in the definition of $\delta$. The latter issue is a consequence of the fact that baryons carry negative strangeness. Ultimately, this is due to the transformation properties of the representations of $SU(3)$ under charge conjugation.
Figure 1: Charge fractions of SNM for $T=\mu_{Q}=\mu_{S}=0$. The solid, black line is the charge fraction $Y_{Q}(\mu_{B})$, the dashed, red line is the strangeness fraction $Y_{S}(\mu_{B})$, and the dot-dashed, blue line that combines by the charge fraction and strangeness contributions to properly account for isospin symmetry.
The parameter $\delta_{I}$ solves other issues with the usual $\delta_{Q}$ parameter in the case where protons are not favored over neutrons, i.e. for vanishing isospin chemical potential. For instance, for matter composed of neutrons and protons only, $Y_{Q}=\frac{n_{p}}{n_{p}+n_{n}}=\frac{1}{2}$ for $\mu_{I}=0$, yielding zero asymmetry $\delta_{Q}=0$, as expected. However, this fails in the presence of other (strange) particles. For instance, in the presence of $\Lambda^{0}$ and $\Sigma^{\pm}$ hyperons, $Y_{Q}=\frac{n_{p}}{n_{p}+n_{n}+n_{\Lambda}}<\frac{1}{2}$, such that $\delta_{Q}>0$, even for $\mu_{I}=0$. If $\delta_{I}$ is used instead, this problem is solved and $\delta_{I}=0$ for $\mu_{I}=0$, even in the presence of neutral and charged hyperons.
Figure 1 illustrates the behavior of $Y_{Q}$ as a function of the baryon chemical potential $\mu_{B}$ for vanishing temperature and $\mu_{Q}=T=0$. The displayed results were obtained for $\mu_{S}=0$ (in weak equilibrium) with the recent CMF++ code Cruz-Camacho et al. (2024) for C2 coupling, which includes the full SU(3) baryon octet and decuplet as well as quarks. We find that the change in $Y_{Q}<1/2$ occurs precisely when $Y_{S}\neq 0$. If we instead consider the combination $Y_{Q}-Y_{S}/2$ at $\mu_{Q}=0$, then this change is absent and its value remains constant at $1/2$, as illustrated by the blue dot-dashed curve in Fig. 1.
Figure 2: The isospin chemical potential $\mu_{I}(n_{B},\delta)$ for positive and negative values of the asymmetry parameter in its original form, $\delta_{Q}$ (top), in our new $\delta_{I}$ form for isospin symmetric matter $\mu_{S}=-1/2\mu_{Q}$ (middle), and in after redefinition, $\delta_{S}$ for weak equilibrium $\mu_{S}=0$ (bottom).
Another way to look at the isospin reflection symmetry is via $\mu_{I}(\delta)$, which should be zero for $\delta=0$. For ANM, we can check whether $\mu_{I}(-\delta)=-\mu_{I}(\delta)$, which is expected from isospin symmetry. In Fig. 2 we show $|\mu_{I}|(\delta)$ for different choices of $\delta_{Q}$ (top) and $\delta_{I}$ (middle,bottom). We find that $\mu_{I}\neq 0$ for $\delta_{Q}=0$ at high $n_{B}$ because of the appearance of strange particles, which implies once again that $\delta_{Q}$ is a poor basis for the expansion. We then consider two cases for the $\delta_{I}$ term comparison:
  • Case 1: Isospin symmetric expansion in $\delta_{I}$. In Eq. (4) we can obtain a reflection in isospin by setting $\mu_{S}=-\mu_{Q}/2$ such that each fixed $n_{B}$, the energy change is same as $I_{z,i}\to-I_{z,i}$.
  • Case 2: Strange break of Isospin Symmetry in $\delta_{I}$. In the case of weak equilibrium $\mu_{S}=0$, isospin reflection is broken by the contribution of a $(1/2)\mathcal{S}\mu_{Q}$ term as shown by Eq. (4).
For case 1 where $\mu_{S}=-\mu_{Q}/2$ we find perfect isospin reflection of $|\mu_{I}|(+\delta_{I})=|\mu_{I}|(-\delta_{I})$ even after the onset of strangeness ($n_{B}\gtrsim 0.57\,fm^{-3}$ for SNM). For case 2 where $\mu_{S}=0$ we find a significant improvement when using $\delta_{I}$ (compared to $\delta_{Q}$) such that there is perfect agreement before strangeness switches on, but after $n_{S}\neq 0$ then we find $|\mu_{I}|(\delta_{I})>|\mu_{I}|(-\delta_{I})$. This relationship makes sense given the contribution of the $(1/2)\mathcal{S}\mu_{Q}$ that contributes positively for $\delta_{I}>0$ ($\mathcal{S}<0$ and $\mu_{Q}<0$) and negatively for $\delta_{I}>0$ ($\mathcal{S}<0$ but $\mu_{Q}>0$).
Before delving into the details of these two cases, we want to introduce the symmetry energy expansion. Originally, the binding energy across a range of nuclei suggested that $\tilde{E}(\delta_{Q})=\tilde{E}(-\delta_{Q})$ Bombaci and Lombardo (1991), which was leveraged to eliminate odd terms in $\delta_{Q}$ and $\delta_{Q}^{3}$. For just protons and neutrons, $\delta_{Q}\to-\delta_{Q}$ symmetry is a good approximation (with certain caveats Roca-Maza et al. (2018)). Theoretically, this parity observation can be understood as a consequence of the isospin reflection symmetry, so we may expect $\tilde{E}(\delta_{I})=\tilde{E}(-\delta_{I})$ even after the onset of strangeness. One expands the energy per nucleon $\tilde{E}/A$ of ANM at some value of $\delta$ around that of SNM, in powers of the asymmetry parameter $\delta$ (here either $\delta_{Q}$ or $\delta_{I}$):
$\frac{\tilde{E}_{ANM}}{A}(n_{B},\delta)$ $=$ $\frac{\tilde{E}_{SNM}}{A}(n_{B})+S(n_{B},\delta),$ (9a)
$S(n_{B},\delta)$ $=$ $\tilde{E}_{sym,1}(n_{B})\,\delta+\tilde{E}_{sym,2}(n_{B})\,\delta^{2}$ (9b)
$+$ $\tilde{E}_{sym,3}(n_{B})\,\delta^{3}+\mathcal{O}(\delta^{4})$
where $S(n_{B})$ is defined as the difference between the ANM and SNM energies and we subtract the nucleon rest mass $m_{np}=(m_{n}+m_{p})/2$ from $\tilde{E}/A=\varepsilon/n_{B}-m_{np}$, where $\varepsilon$ is the energy density.
Figure 3: Energy per baryon number vs $\delta$ at a fixed baryon density $n_{B}=0.99\;[fm^{-3}]$. The original $\delta_{Q}$ finds that the ground state of nuclear matter is at a finite $\delta_{Q}\sim 0.2$ whereas our new $\delta_{I}$ has the minimum coincide with $\delta_{I}=0$. Thus, all linear terms should disappear from the symmetry energy expansion. The binding energy is symmetric across $\delta_{I}$.
Case 1: Strangeness isospin symmetry We now consider the case of isospin symmetric matter that has a reflection symmetry across the SNM axis. To do so, we take $\mu_{S}=-1/2\mu_{Q}$ within CMF and study the behavior of $\tilde{E}/A$ across either $\delta_{Q}$ or $\delta_{I}$, as shown in Fig. 3 at $n_{B}=0.99\,fm^{-3}$ that is well above the onset of strangeness. We find that $\delta_{Q}$ leads to a significant problem in that the ground state of nuclear matter shifts to a finite $\delta_{Q}\sim 0.2$ value such that one would obtain linear terms in the symmetry energy expansion i.e. $\tilde{E}_{sym,1}\neq 0$. However, if we instead use $\delta_{I}$ it is clear that we have isospin reflection symmetry across the $\delta_{I}=0$ axis and that an expansion up to $\delta_{I}^{2}$ leads to a nearly perfect reproduction of the thermodynamics.
Thus, for the isospin reflection case for strangeness i.e. when $\mu_{S}=1/2\mu_{Q}$ our symmetry energy expansion can be written as:
$\frac{\tilde{E}_{ANM}}{A}(n_{B},\delta_{I})\big{|}_{\mu_{S}=\frac{\mu_{Q}}{2}}=\frac{\tilde{E}_{SNM}}{A}(n_{B})+\tilde{E}_{sym,2}(n_{B})\,\delta_{I}^{2}$ (10)
where in the Supplemental Materials we demonstrate a surprisingly accurate description of $\tilde{E}_{sym,2}(n_{B})$ expanding around $n_{sat}$ up to $\mathcal{O}(n_{B}^{2})$.
Case 2: Strange Breaking of Isospin Symmetry In the case of $\beta$-equilibrium (otherwise known as weak equilibrium), both decays are in equilibrium:
$d$$\rightarrow$ $u+e^{-}+\bar{\nu}_{e}$ (11)
$s$$\rightarrow$ $u+e^{-}+\bar{\nu}_{e}$ (12)
such that the chemical potentials must hold:
$\tilde{\mu}_{d}$ $=$ $\tilde{\mu}_{u}+\tilde{\mu}_{e}$ (13)
$\tilde{\mu}_{s}$ $=$ $\tilde{\mu}_{u}+\tilde{\mu}_{e}$ (14)
$\therefore\tilde{\mu}_{d}$ $=$ $\tilde{\mu}_{s}$ (15)
$\therefore\mu_{S}$ $=$ 00 0 (16)
which we anticipate within old, isolated neutron stars. Here, we do not anticipate perfect isospin reflection symmetry, but we expect that the choice of $\delta_{I}$ will still significantly improve the symmetry energy expansion.
In Fig. 4 we show the energy per baryon, $\varepsilon/n_{B}(n_{B})$, for SNM and $\delta_{I}=\pm 0.3$. We find a perfect symmetry under $\delta_{I}\to-\delta_{I}$ at low $n_{B}$ and a slight breaking of this symmetry above $n_{B}\gtrsim 0.57\,fm^{-3}$, but we see an even more pronounced breaking for $n_{B}\gtrsim 0.6\,fm^{-3}$, with $\varepsilon/n_{B}(\delta_{I})>\varepsilon/n_{B}(-\delta_{I})$, which is consistent with the effect in $|\mu_{I}|(\pm\delta_{I})$.
Figure 4: Energy density over baryon density vs baryon density for symmetric nuclear matter $\delta_{I}=0$, and demonstrating reflection symmetry across the $\delta_{I}$ axis by comparing $\delta_{I}=0.3$ vs $\delta_{I}=-0.3$. At low $n_{B}$, $Y_{S}=0$ such that we see a reflection symmetry at finite $\delta_{I}$ but at
Figure 5: Energy per baryon number vs $\delta_{I}$ at a fixed baryon density $n_{B}=0.76\;[fm^{-3}]$. The original $\delta_{Q}$ finds that the ground state of nuclear matter is at a finite $\delta_{Q}\sim 0.1$ whereas our new $\delta_{I}$ has the minimum coincide with $\delta_{I}=0$. Thus, all linear terms should disappear from the symmetry energy expansion. We can see that the binding energy is asymmetric across $\delta_{I}$ such that a skewness term ($\delta_{I}^{3}$) provides a nearly perfect fit. The green cross marks the onset of strangeness. On the left, $Y_{S}=0$, while on the right, $Y_{S}\neq 0$.
In Fig. 5 we show $\tilde{E}/A(n_{B})$ and fit a polynomial in $\delta_{I}$ or $\delta_{Q}$ to the data. The isospin reflection symmetry is broken due to finite $Y_{S}$, as expected. Thus, when $Y_{S}\neq 0$:
  1. 1.
    Using $\delta_{Q}$ shifts the ground state of nuclear matter to a finite $\delta_{Q}\sim 0.03$, breaking linear symmetry.
  2. 2.
    Using $\delta_{I}$ shifts the ground state of nuclear matter back to $\delta_{I}=0$, restorying linear symmetry.
  3. 3.
    Weak equilibrium ($\mu_{S}=0$) leads to a skewness across $\delta_{I}$ such that $E_{\textrm{sym},3}\neq 0$ in Eq. (9b).
Figure 6: Contour plot of constant strangeness fraction $Y_{S}$ projected into the plane of baryon density vs. $\delta_{I}$, with isospin symmetric system (top), and system in weak equilibrium (bottom).
Phase diagram of $Y_{S}$ One of the most crucial differences to our two cases listed above is the phase diagram of $Y_{S}(n_{B},\delta_{I})$. In Case 2 of weak equilibrium, $\mu_{S}=0$ such that $\mu_{S}^{(B,I_{z})}=\frac{1}{2}\mu_{Q}$. Positive $\delta_{I}$ leads to $\mu_{I}<0$ (recall in this limit $\mu_{Q}=\mu_{I}$) such that $\mu_{S}^{(B,I_{z})}<0$ and since hyperons carry negative strangeness it is easier to produce strange particles. In contrast, $\delta_{I}<0$ suppresses strangeness. Once $Y_{S}\neq 0$, this leads to asymmetries under $\delta_{I}\to-\delta_{I}$. In contrast, in Case 1 where $\mu_{S}^{(B,I_{z})}=0$ then there is an equal probability to produce strangeness in both the ±$\pm\mu_{I}$ directions such that $Y_{S}$ is symmetric around the $\delta_{I}$ axis.
In Fig. 6 we can validate these arguments where Case 1 is shown (top) to have $Y_{S}$ that is, indeed, symmetric around the $\delta_{S}$ axis. In contrast we find that $Y_{S}$ has a clear skewness across the $\delta_{S}$ axis because it is easier to produce strangeness for $\delta_{S}>0$ whereas strangeness is suppressed for $\delta<0$. This skewness in strangeness for the weak equilibrium case is precisely what leads to the appearance the symmetry breaking strangeness term $E_{\textrm{sym},3}\neq 0$.
Conclusions and Outlook In this paper, we found that the typical isospin asymmetry coefficient $\delta_{Q}=1-2Y_{Q}$ does not obtain the ground state of nuclear matter, when strangeness is finite. Thus, we have rewritten the isospin asymmetry coefficient $\delta_{I}$ to be inclusive of strangeness, which can obtain the ground state of nuclear matter for symmetric nuclear matter both with and without strangeness. We then consider two cases: one that preserved isospin symmetry in the presence of strangeness $\mu_{S}=-1/2\mu_{Q}$ and the other case associated with old, cold neutron stars of weak equilibrium $\mu_{S}=0$ that breaks isospin symmetry.
Our results have led to two very interesting findings that can pave the way for future work. The first finding is that $Y_{S}\neq 0$ leads to a skewness around the ground state of nuclear matter in the case of weak equilibrium (given a specific $n_{B}$). In other words, strangeness implies isospin symmetry breaking across the SNM axis when $\mu_{S}=0$. It is likely possible to mathematically connect the isospin symmetric system of $\mu_{S}=-1/2\mu_{Q}$ to that of weak equilibrium, which we leave for a future work. However, we believe that this strangeness breaking of isospin symmetry may help to explain some of the interesting phase transitions that appear at $T=0$ across finite $\mu_{S}$ and $\mu_{Q}$ in CMF++ Cruz-Camacho et al. (2024).
The second implication of our findings is that new methods are needed to extract symmetry energy coefficients in the presence of strangeness from experimental data, chiral effective field theory, and other theoretical models. One potential possibility is creating systems that include finite strangeness. There are two possible types of experiments that we envision that could provide information about strangeness. The first would be fixed-target collisions of kaon beams (see e.g. NA6 (2022) other experiments with kaon beams) with a nucleus and the second would be collisions of hypernuclei. Similarly, it may be possible to obtain information about these coefficients in chiral effective field theory if strange interactions were included (see e.g. Haidenbauer and Meissner (2010); Li et al. (2018); Song et al. (2018); Haidenbauer et al. (2016)). Finally, varying the large parameter space of CMF++ (and other similar EOS models that include hyperons) while constraining it to reproduce saturation properties and neutron star properties, can help to extract the allowable range of these symmetry coefficients.
Acknowledgments This research was partly supported by the US-DOE Nuclear Science Grant No. DE-SC0023861 and by the National Science Foundation (NSF) within the framework of the MUSES collaboration, under grant number OAC-2103680. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. We also acknowledge support from the Illinois Campus Cluster, a computing resource that is operated by the Illinois Campus Cluster Program (ICCP) in conjunction with the National Center for Supercomputing Applications (NCSA), which is supported by funds from the University of Illinois at Urbana-Champaign.